OpenStudy (anonymous):

Gas laws help Please!! What volume of nitrogen at 215°C and 715 mm Hg would be required to react with 0.100 mol of hydrogen to produce ammonia? N2(g) +3H2 (g) --> 2NH3(g)

OpenStudy (anonymous):

I have tried PV=nRT; the question has all the variables required to use Avogadro's law. The answer I get is 4.26L but it says that's wrong.

OpenStudy (anonymous):

Okay I got the correct answer which is 1.42L. I didn't pay attention to the stoich; how can you properly apply stoich to this though? I just multiplied by 1/3, but isn't there a proper way to put this into dimensional analysis?

OpenStudy (jfraser):

To react with 0.100mol of hydrogen, you need to use 1/3 that much nitrogen, that's what the balanced reaction does.\[0.100mol \space H_2 * \frac{1mol \space N_2}{3mol \space H_2}\]

OpenStudy (jfraser):

then plug that value into the IGL with the rest of the given info to find the volume

OpenStudy (anonymous):

Okay. I've seen IGL being used in DA though.. As in, the whole problem was done using DA, that seems quite a bit easier than trying to use algebra, do you know how that works?

OpenStudy (jfraser):

since the IGL is\[P*V = n*R*T\], then solving for volume gets you\[V = \frac{n*R*T}{P}\]

OpenStudy (jfraser):

using DA, and starting with the moles of gas looks like this

OpenStudy (jfraser):

\[0.100mol \space H_2 * \frac{1mol \space N_2}{3mol \space H_2}*\frac{0.0821L*atm}{mol*K}*\frac{488K}{1}*\frac{1}{0.941atm}\]